OpenStudy (anonymous):

What is the slope of the line –6y + 2x = 18 in simplest form?

OpenStudy (anonymous):

Change it to slope intercept form. It's y = mx + b. -6y + 2x = 18 First we subtract 2 to both sides: -6y = -2x + 8 Now we divide -6 to both sides, what's -2 / -6?

OpenStudy (adamk):

To find the slope, let's first put the equation in slope-intercept form, or y = mx + b. So that would be y = \[\frac{ 1 }{ 3 }x\] -3

OpenStudy (adamk):

Then, to find the slope, we just take "m" of y = mx + b. So the coefficient of "x" in your new equation.

OpenStudy (anonymous):

Actually, we should simplify 2/6, can you do that?

OpenStudy (anonymous):

3 @iGreen.

OpenStudy (adamk):

@edennoyes I think you meant 1/3

OpenStudy (adamk):

;)

OpenStudy (anonymous):

or 0.3 @iGreen.

OpenStudy (anonymous):

Yeah, and that would be our slope.

OpenStudy (anonymous):

Slope is 1/3 or 0.3

OpenStudy (anonymous):

\[–6y + 2x = 18\\-6y=-2x+18\\y=\frac{2}{6}x-\frac{18}{6}\\y=\frac{1}{3}-3\\\boxed{m=\dfrac{1}{3}}\]

OpenStudy (anonymous):

ohhhh! thanks! @iGreen. @AdamK